x^y+y^x=1500, then find the value of x and y?

Grade:upto college level

2 Answers

PERUMAL
10 years ago
x^y+y^x=1500 1499^1+1^1499=1500 x=1499, and y=1
PERUMAL
10 years ago
x^y+y^x=1500 1499^1+1^1499=1500 x=1499, and y=1 this problem is prove.